Measurement and Error Analysis Data Sheet Table 1. Measure the dimensions of the cylinder using the vernier caliper Zero correction:Indication error O-, Unit: Trial Outer diameter Inner diameter Height 2 ?1008-26.oel n? toon :n.07| so t o.og: O.? 4 5, , 35.471 Di Measure the diameter of the ball using the micrometer caliper 0.004 Table 2. Zero correction:Indication error:Unit: m Trial 0-e2 Diameter D-3.2 4-24 4.24 14.26 A-75 4. 26 -1 X16 o-ol 4 0.01
media%2Fafb%2Fafbb503b-298c-423f-b6e7-bc


media%2Ffd9%2Ffd96889b-3ccd-44c7-bb5b-e5
media%2F06a%2F06a827d8-709d-49e3-863f-60
Answer #1

Given

\bar D_1=35.472mm, \bar D_2=22.038mm , \bar H=50.09mm

Assuming diameter and height, both are measured with vernier calliper.

So we get

\Delta D_1= \Delta D_2 = indication error of the vernier calliper= 0.02mm

\Delta H = indication error of the vernier calliper= 0.02mm

Volume of the cylinder given by

V=\frac{\pi}{4}(\bar D_1^2-\bar D_2^2)\bar H

  =\frac{\pi}{4}(35.472^2-22.038^2)\times 50.09= 30394 mm^3

\Delta V=\sqrt{(\frac{\pi}{2}\bar D_1\bar H \Delta D_1)^2+(\frac{\pi}{2}\bar D_2\bar H \Delta D_2)^2+[\frac{\pi}{4}(\bar D_1^2-\bar D_2^2) \Delta H ]^2}

=\sqrt{55.82^2+34.68^2+12.14^2}=66.827 (mm^3)

Given diameter of the ball \bar D = 14.25 mm

\Delta D = indication error of the vernier calliper= 0.004mm

So, Volume of the ball is

\bar V= \frac{\pi \bar D^3}{6}=1515 mm^3

\Delta V= \frac{3 \Delta D}{D} \bar V=\frac{3 \times 0.004}{14.25} \times 1515=1.275 mm^3

3. Given Mass of the cylinder M= 86.8 gm

Density of the cylinder, \bar \rho=\frac{M}{\bar V}=\frac{86.8}{30394}(gm/mm^3)= 2855 (gm/cm^3)

Taking logarithm of the above equation

ln \bar \rho=ln M- ln \bar V

and differentaiting both side we get

\frac{\Delta \bar \rho}{ \bar \rho}=\frac{\Delta M}{M}-\frac{\Delta \bar V}{\bar V}

since errors are always add up we can write

\frac{\Delta \bar \rho}{ \bar \rho}=\frac{\Delta M}{M}+\frac{\Delta \bar V}{\bar V}

Or

\Delta \bar \rho=[\frac{\Delta M}{M}+\frac{\Delta \bar V}{\bar V}]\bar \rho

\Delta M = Let indication error of the vernier calliper= 0.01mg( As it is not given)

So

\Delta \bar \rho=[\frac{0.01}{86.8}+\frac{1.275}{30394}] \times 2855= 0.448 (gm/cm^3)

So

\rho= \bar \rho \pm \Delta \bar \rho

\rho=2855 \pm 0.448 (gm/cm^3)